# Bridging Developer Needs and Feasible Features for AI Assistants in IDEs

### v2

@@ III-B FINDINGS
In Figure 1, we present our categorization of developer needs regarding Human-AI Experience within IDEs, based on a thematic analysis of 35 interviews. It includes five overarching topics: Technology Improvements, Simplifying Programming Tasks, Technology Interaction, Technology Customization, and Simplifying Skill Building.
@@
Participants stressed the need for more reliable AI systems that reflect the latest advancements of the field. Participants also raised concerns about the exhaustion of training data, the perpetuation of outdated or erroneous information, and the plateauing progress of AI, all of which could undermine the effectiveness and accuracy of tooling. Participants also noted current shortcomings in output quality and emphasized the importance of thoroughly reviewing and validating AI-generated code to avoid overreliance.
@@
Privacy and security were also identified as critical requirements for AI adoption. Participants favored on-premise or local AI solutions to ensure data confidentiality and comply with organization-specific or national restrictions. They stressed the importance of preventing vulnerabilities and maintaining the integrity of AI-generated code, including making such code clearly identifiable to facilitate inspection and verification.
“Regulations restrict us from using AI when it involves sending data to the cloud. My company is particularly interested in on-premise language models to maintain control and keep data within the country or even company.”A77
Context awareness was seen as essential for generating relevant and accurate code. The participants highlighted the need for mechanisms to manually set, view, and edit the context provided to the AI assistant. Some expressed frustration over the lack of control in current systems and noted that greater transparency in how context is constructed could also help mitigate privacy concerns.
The participants also called for better alignment of AI tools with their existing development practices. They praised models that generate suggestions consistent with the project structure and coding conventions, such as recommending appropriate libraries or learning from examples. At the same time, they noted issues with overly verbose, redundant, or mismatched output that diminished the value of AI-generated assistance.
Proactive features were seen as an opportunity to shift AI from a reactive assistant to a more autonomous collaborator. Participants envisioned systems that monitor ongoing work, detect issues, and make targeted improvements, such as flagging performance bottlenecks or managing dependencies, as long as these actions remain in the trusted IDE environment and do not involve external model training.
Accessibility was another recurring concern. The interviewees pointed to the need for tools that are affordable and adaptable across infrastructural, legal, and experiential boundaries. Suggestions included natural language programming, universal language support, and modular pricing models that lower the barrier to entry, particularly in regions with limited connectivity or restrictive regulations.
The participants also emphasized the importance of ethical safeguards. They called for clear rules and restrictions to guide responsible AI use and suggested that automated mechanisms may be necessary to ensure compliance and accountability,particularly as AI becomes more integrated into daily development workflows.
@@
The participants emphasized the potential of AI to reduce manual effort by supporting routine and repetitive activities across the software development lifecycle. By automating low-level tasks during ideation, implementation, testing, maintenance, and optimization, developers could shift focus to more complex and creative work. Many participants envisioned AI evolving into an active collaborator capable of supporting ideation, implementation, testing, maintenance, and optimization.
@@
A common expectation was for AI to handle well-defined repetitive work, such as template-based coding, documentation generation, and routine debugging. This division of labor was seen as essential for maintaining developer productivity without increasing cognitive load.
The interviewees also emphasized the importance of AI in supporting codebase navigation. In large or unfamiliar projects, features such as database analytics, diagram generation, and in-context code explanations were considered particularly valuable. These capabilities were viewed as instrumental in reducing ramp-up time and enabling a more seamless understanding of system architecture.
@@
The participants also expressed interest in AI tools that integrate IDE with external systems, such as CI/CD pipelines, version control platforms, and support ticketing systems. The ability to access and coordinate AI assistance across these environments was seen as a way to increase traceability and streamline workflows spanning multiple tools.
@@
The interviewees emphasized the importance of smooth internal system integration, where AI tools interact efficiently with other features within the IDE. Integrating AI into the IDE, especially through embedded chat functionality, was considered particularly beneficial, reducing context switching and, therefore, avoiding unnecessary interruptions.
“I don’t need to switch to a different tab or window. I don’t have to go through that whole process, so that’s definitely very helpful. It saves time and makes the process more efficient.”C122
The participants expressed a desire for other types of uninterrupted experiences that allow them to interact with AI in flexible ways, including via voice commands or wearables such as Neuralink. They also showed interest in features like inter-AI interaction, in-line code explanations, and customizable key mappings to further streamline workflows.
At the same time, several challenges were reported. These included slow response times, frequent logouts, and limitations in chat functionality, such as copy/paste issues or dead-ends in conversations where the AI failed to respond. Participants also noted struggles with reproducibility in code completion, UI inconsistencies, and interference between suggestions from different completion providers. Such issues were seen as disruptive to overall workflow continuity.
  • *Note: Symbols represent the participant groups: Adopters (⚫), Churners (★), and Non-Users (▲). Presence of a symbol indicates that group members mentioned the corresponding theme in the associated interview section.
Figure 2: Comparative Analysis of Thematic Group Mentions Across User Types.
@@
The interviewees stressed the importance of adjustable AI systems that offer a high degree of personalization and flexibility, allowing users to retain control as decision-makers. Key features included configurable AI, the ability to choose between different models, and the option to customize AI actions based on specific tasks or workflows. Participants also expressed a desire for task-specific setups that allow AI to adapt to unique user preferences and working styles.
To support workflow optimization, participants were also interested in tools that could help manage distractions, track time, and summarize daily activities within their development ecosystem. The ability to toggle time tracking, manage task tickets, and maintain a personal prompt library was highlighted as essential for staying organized and improving productivity.
@@
Answer to RQ1. Developers emphasized the need for accessible, proactive, reliable, secure, and context-aware technology that adapts to their workflows and remains ethical and up-to-date. They seek AI tools that support various aspects of development, from code generation and testing to codebase navigation, focusing on non-interruptive, customizable, and learning-supportive interfaces. Moreover, developers prioritize smooth internal integration of AI with existing systems, emphasizing the need for intuitive user education and skill development to maximize AI’s potential within their workflows.
Comparative Analysis of Codes Among Groups
We present a comparative analysis of the Thematic groups among Adopters, Churners, and Non-Users in Figure 2, focusing on participants’ success stories (or anticipated advantages for Non-Users), challenges (or barriers for Non-Users), and needs in the context of in-IDE HAX. Lastly, we explore the similarities and differences at the Codes’ level.
  • Similarities Across All Groups:
    Concerns about AI displacing jobs were widespread, reflecting apprehension about job security and the need for AI tools to complement rather than replace human capabilities. Participants expressed a general interest in AI’s role in code generation, explanation, and evaluation, recognizing its potential to simplify coding tasks. We identified the lack of sufficient project-wide context awareness in AI tools as a common challenge, with participants calling for improvements in this area to boost AI’s effectiveness.
    Ethical and privacy concerns, such as plagiarism and data protection, were also frequently mentioned, emphasizing the need to address these issues to build trust. The participants also noted the risk of overreliance on AI, highlighting the importance of balancing AI assistance with personal skill development. Despite these concerns, there was a shared sense of optimism, with many participants acknowledging the ongoing improvement of AI tools over time.
  • Similarities Between Adopters and Churners:
    Both groups faced challenges with AI integration, such as code completion clashes and frequent logouts, confirming that such technical difficulties negatively affect user retention and satisfaction. There was a shared interest in AI functionalities that proactively assist with debugging, optimization, and documentation, stressing a demand for more advanced AI capabilities. While some Adopters and Churners reported similar challenges when using AI assistants, Adopters were more likely to persist through issues, experiment with workarounds, and continue using the tools despite their limitations. This contrast suggests that factors beyond the AI’s technical performance influence user retention. Further research is needed to understand what shapes developers’ willingness to tolerate imperfections and continue engaging with AI tools over time.
  • Differences Between Adopters and Churners:
    Adopters demonstrated a more positive outlook, frequently expressing optimism and enthusiasm. They also reported interest in advanced features like Inter-AI Interaction and Cross-Model Prompting, reflecting a deeper integration of AI into their workflows. Additionally, Adopters showed a clear preference for customization and personalization of AI tools to fit their specific needs.
    In contrast, Churners often exhibited skepticism and dissatisfaction. They emphasized security and privacy concerns more, favoring on-premise AI solutions due to trust issues with cloud-based services. Churners also highlighted technical limitations such as slow response times, outdated data, and limited creativity from language models, which contributed to their discontinuation of AI usage. It is worth noting that several Churners were forced to stop their AI usage due to company or country restrictions. They mentioned being overall satisfied with their experience and emphasized that more secure and on-premise solutions could get their companies to reconsider the restrictions.
  • Unique Themes Among Non-Users: Despite mostly being aware of AI capabilities, Non-Users had not adopted these tools, indicating potential barriers to initial adoption. A strong focus on personal skill development emerged among Non-Users, suggesting a preference for traditional learning methods over AI assistance. Moreover, Non-Users expressed unique ethical concerns about the environmental impact and ownership issues, highlighting the influence of broader societal implications on their decision not to engage with AI technology.
Answer to RQ2. Adopters show enthusiasm for advanced AI features and customization, suggesting that further innovation can strengthen their engagement. Churners, however, highlight technical issues and trust concerns. This indicates the need for enhanced reliability and privacy-preserving measures. Non-Users emphasize personal skill development and ethical concerns and highlight a barrier to adoption that can be mitigated with better education and clearer ethical guidelines.
IV ESTIMATING FEATURE FEASIBILITY VIA PREDICTION MARKET
Table I: Prediction accuracy and market activity across tools for proposed features.
Feature Tool Resolved Acc.? Prob. Volume Liq.
@@ IV ESTIMATING FEATURE FEASIBILITY VIA PREDICTION MARKET
Project-wide Context Awareness Any Tool 82% 440 290
@@
Proactive Documentation Any Tool 33% 878 310
@@
Proactive Automation Any Tool 24% 590 290
JetBrains IDE 13% 440 270
Annotating AI-Generated Code Any Tool 18% 733 290
JetBrains IDE 18% 640 270
Business-context Awareness Any Tool 22% 272 150
@@
Daily Ecosystem Summarizer Any Tool 15% 295 190
JetBrains IDE 14% 510 250
AI Real-Time Refactoring Any Tool 19% 950 370
@@
Proactive Optimization Any Tool 19% 650 310
@@
Proactive Debugging Any Tool 6% 750 210
@@
Brain-computer Interface Coding Any Tool 4% 1241 490
@@
  • *Note: Each row shows a market prediction about a feature launch. Resolved shows if the feature was launched (✓) or not (✗). Acc.? indicates if the prediction was correct. Prob. is the final market probability. Volume and Liq. reflect participant engagement.
In the first phase of the study, described in Section III, we explored the needs of different groups of users in regards to AI. In the subsequent phase, we investigated how the market responds to these needs in a short-term period. We were also interested in whether the tech creators’ community can successfully predict which of the needs are likely to be met. Therefore, we conducted a prediction market “tournament” on JetBrains’ internal platform.
Prediction markets are exchange-traded markets where people’s bets, with payoffs depending on a future event, can be used as a measure of the common beliefs. The prevailing price of the bet represents the collective belief in the probability of that event. This effect follows from the principle that aggregating many independent bets yields accurate forecasts, the so-called wisdom of the crowd [10]. Previously, prediction markets have been successfully used to forecast a broad range of outcomes from presidential votes [29], to replicability of scientific research [30], and the future of real products on the market [31]. In our research, through prediction market bets, we examined which user needs and feature directions tool builders recognize and consider relevant. We were also interested in testing whether these bets reflect an accurate intuition about how the field is likely to evolve, measured by whether the predicted features are eventually implemented or not. Consequently, our research question for Phase 2 is
  • RQ3. To what extent do practitioners recognize and correctly anticipate which AI features inspired by user needs are likely to be implemented, as reflected in their prediction market activity?
@@
All materials used for data gathering at this phase, as well as the full demography table and a summarized report of the prediction market, are in Supplementary materials [28].
IV-A METHOD
IV-A1 Materials

### v3

@@ 3.2.1. Categorization of Developer Needs
Figure 1, presents our categorization of developer needs regarding Human-AI Experience within IDEs, based on a thematic analysis of 35 interviews. It includes five overarching topics: Technology Improvements, Simplifying Programming Tasks, Technology Interaction, Technology Customization, and Simplifying Skill Building.
Refer to caption A two-column table showing 25 thematic groups organized by topic area, with columns indicating which participant groups (Adopters, Churners, Non-Users) mentioned each theme as Successes, Challenges, or Needs. The left column contains technology-related themes including Accessible Technology, Proactive Technology, Reliable Technology, Up-to-date Technology, Ethical Technology, Private Technology, Secure Technology, Aligned Technology, Context-aware Technology, Non-Interruptive Technology, and Internal System Integration. The right column contains task and interaction themes including AI Support in SDLC, Navigating the Codebase, AI in Ideation, AI in Implementation, AI in Testing, AI in Maintenance, Code Optimization, External System Integration, Adjustable Technology, Optimizing Workflow, User Education, and Learning Support. Symbols indicate group mentions: filled circles for Adopters, stars for Churners, and triangles for Non-Users. Themes are color-coded by category: orange for Technology Improvement, teal for Interaction, blue for Customization, purple for Programming Tasks, and green for Simplifying Skill Building. The table reveals that Adopters reported more Successes across themes, while Churners and Non-Users more frequently mentioned Challenges and Needs, particularly in areas like Proactive Technology, Context-aware Technology, and AI in Maintenance.
Note: Symbols represent the participant groups: Adopters (⚫), Churners (★), and Non-Users (▲). The presence of a symbol indicates that group members mentioned the corresponding theme in the associated interview section.
Figure 2. Comparative Analysis of Thematic Group Mentions Across User Types.
@@
Participants identified several requirements for improvements of AI tooling for coding, beginning with accessibility across infrastructural, legal, and experiential boundaries. They proposed specific solutions, including natural language programming interfaces, universal language support, and modular pricing models, to reduce barriers in regions with limited connectivity or restrictive regulations. Beyond basic accessibility, participants envisioned AI systems transitioning from reactive assistants to proactive collaborators. This shift would enable AI to identify potential issues during development:
Such proactive capabilities, however, require fundamental improvements in system reliability and integration of the latest advancements in the field. Participants emphasized that unreliable systems create additional verification overhead:
@@
These technical improvements must be accompanied by robust ethical frameworks. Participants recognized AI’s potential and the corresponding responsibility this creates:
@@
Ethical considerations extend to privacy and security requirements. Participants strongly favored on-premise or local AI solutions to ensure data confidentiality and comply with organizational or national restrictions. They emphasized the need for vulnerability prevention and maintaining code integrity, including mechanisms to clearly identify AI-generated code for inspection and verification. Finally, participants stressed that all improvements must support better alignment with existing development practices. Context awareness emerged as essential for generating relevant and accurate code. They specifically requested mechanisms to manually set, view, and edit the context provided to AI assistants. Current systems’ lack of transparency in context construction creates frustration and undermines user confidence in AI-generated outputs.
@@
Participants identified AI’s primary value in automating routine and repetitive activities across the software development lifecycle (SDLC). This automation would enable developers to redirect effort from low-level tasks toward more complex and creative work spanning ideation, implementation, testing, maintenance, and optimization phases. This vision extends beyond simple task automation to AI functioning as an intelligent collaborator:
@@
A critical component of this collaborative relationship involves AI-supported codebase navigation. Participants particularly valued AI’s potential to help newcomers understand existing systems and adhere to established standards:
@@
Participants also envisioned AI tools that seamlessly connect IDEs with external environments, including Continuous Integration and Continuous Deployment (CI/CD) pipelines, version control platforms, and support ticketing systems.
@@
Participants identified seamless internal system integration as fundamental to an effective AI tool, emphasizing the need for AI to interact efficiently with existing IDE features. Beyond that, participants sought uninterrupted experiences that enable flexible AI interaction modes throughout their development workflows. However, current implementations frequently fall short of this due to technical limitations that disrupt workflow continuity. These disruptions manifest in several forms: slow response times that break the natural rhythm of development, frequent authentication failures that force repeated logins, and functional limitations including copy/paste failures and conversation dead-ends where AI systems become unresponsive. Such technical barriers undermine the seamless integration that participants consider essential for productive AI-assisted development.
@@
The interviewees stressed the importance of adjustable AI systems that offer a high degree of personalization and flexibility, allowing users to retain control as decision-makers. Key features included configurable AI, the ability to choose between different models, and the option to customize AI actions based on specific tasks or workflows. Participants also expressed a desire for task-specific setups that allow AI to adapt to unique user preferences and working styles. To support workflow optimization, participants were also interested in tools that could help manage distractions, track time, and summarize daily activities within their development ecosystem. The ability to toggle time tracking, manage task tickets, and maintain a personal prompt library was highlighted as essential for staying organized and improving productivity.
@@
[…]
3.2.2. Comparison of Needs’ Themes Among Groups
We present a comparative analysis of the Thematic groups among Adopters, Churners, and Non-Users in Figure 2, focusing on participants’ success stories (or anticipated advantages for Non-Users), challenges (or barriers for Non-Users), and needs in the context of in-IDE HAX. The presence of a symbol indicates that at least one participant within that user group mentioned the corresponding theme during their interview. This binary presence/absence approach was chosen to capture the full spectrum of experiences and concerns across adoption stages, as even singular mentions may represent important but underexplored aspects of the user experience that warrant consideration in feature development.
The Context-aware Technology thematic group is uniquely mentioned by all three groups of participants as a success, a challenge, and a need, indicating the most ambiguous relationship to this aspect of technology. Privacy and User Education emerge as universal concerns and points of potential and urgent improvement, with all three groups reporting challenges and needs in this regard. All three groups of participants mention Learning support and AI Support in various stages of SDLC as success and need, signaling that the evolution of AI tooling is sufficient but still relevant in these areas. Notably, AI in the Ideation theme is not mentioned by any of the groups as a need, signaling that people are less likely to seek AI support at this stage of SDLC. On the other hand, participants mention Internal and External System integration only as a need, implying that this is a valuable attention point for tool builders.
It is interesting to note that participants usually mention in challenges and needs Technology Improvements-related themes, while Simplifying Programming Tasks-related themes are usually mentioned in successes and needs. This trend might be interpreted as a direction for tool builders to improve technology, making it more advanced, as opposed to widening applications, which are already successful enough.
3.2.3. Comparison of Needs’ Codes Among Groups
Further, we explore the similarities and differences of participants’ 78 specific needs at the Codes’ level, represented in Table 1.
Table 1. User group needs across topics.
Topic Code of Need A C N
Accessible Technology Polyglot AI
Separate feature purchase
Reasonable pricing
Proactive Technology Proactive code generation
Proactive documentation
Proactive optimization
Proactive debugging
AI performing actions
File management
Library management
Reliable Technology Providing sources
Output clarifications
Higher quality
Up-To-Date Technology Sota tooling
Ethical Technology Ethical technology
Private Technology Privacy
On-premise AI
Local AI
Secure Technology Annotating ai-code
Security
Aligned Technology Library suggestions
Generation by example
Context-Aware Technology Contextual awareness
Project-wide context awareness
Business-context awareness
Manually set context
Viewable context
High level understanding
AI Support In SDLC Productivity boost
Universal programming assistance
Navigating The Codebase Onboarding support
Database analytics
Diagram generation
Code explanations
AI In Implementation Project set-up assistance
Full stack generation
Ui generation
Multi-file generation
Sketch2code
AI In Testing Coverage report
System testing
Ui testing
Error handling
AI In Maintenance Repository management
Database merging
Dependency management
Guideline enforcement
Code Optimization Optimization
Code review support
Code evaluation
External System Integration Commit management
AI in ci/cd pipelines
Out-ide AI access
External system integration
Non Interruptive Technology Voice interaction
Stable AI functionality
Neuralink
Suggestion preview
Internal System Integration In-line code explanation
Key mappings
AI in code with me
Anywhere-in-ide AI access
Internal system integration
Adjustable Technology Personalization
Configurable AI
Choosing models
AI action customization
Optimizing Workflow Toggleable time tracking
Daily ecosystem summarizer
Distraction management
Ticket management
Automation of repetitive tasks
Proactive workflow recommendations
User Education Prompting support
Shareable prompt library
User education
Learning Support Step-by-step output revealing
Learning support
Note: A = Adopter, C = Churner, N = Non-User
All three groups expressed needs for comprehensive development support by AI (universal programming assistance, higher quality output, productivity enhancement, multi-file generation, system testing), enhanced tools’ context awareness (project-wide and business-context awareness, manually set context), infrastructure flexibility (on-premise AI, CI/CD pipeline integration, external system integration), proactive AI capabilities (code generation, optimization, autonomous actions), and improved user experience (suggestion preview, configurable AI, task automation, learning support). The convergence on these needs suggests they represent core requirements for AI-powered development environments
Adopters demonstrated the broadest range of needs, expressing requirements across all identified categories. This comprehensive need profile reflects their active engagement with AI tools and suggests that continued innovation in advanced features will strengthen retention among this group. Drawing from their relatively extensive AI experience, uniquely identified needs are:
  • Advanced tooling and management features (SOTA tooling, File management, Guideline enforcement, Code explanations),
  • Enhanced alignment and reliability (Personalization, AI action customization, Library suggestions, Generation by example, Providing sources, Output clarifications),
  • Deeper integration capabilities (Commit management, out-IDE AI access, AI in Code With Me, Neuralink),
  • Workflow and learning optimization (Daily ecosystem summarizer, Step-by-Step Output Revealing, Shareable Prompt Library).
Churners expressed needs that align substantially with Adopters, sharing 36 common requirements across most topics. However, Churners uniquely emphasized:
  • System stability (Stable AI functionality),
  • Specific conceptual technical capabilities (Polyglot AI, High-level understanding),
  • Infrastructure management (Library management, Repository management),
  • Specialized interface features (In-line Explanation, Sketch2Code).
This pattern suggests that while Churners share Adopters’ vision for comprehensive AI for code, their focus on stability and conceptual depth indicates that current implementations may lack the robustness and sophistication necessary to sustain engagement. Therefore, implementation quality rather than feature availability drives disengagement decisions.
Non-Users expressed needs in specific areas that represent potential adoption catalysts. Their unique requirements centered on foundational concerns including:
  • Ethical technology, privacy protection, and security,
  • Specialized functionality (Database merging, Key mappings),
  • Productivity control features (Toggleable time tracking, Distraction management).
This pattern suggests that Non-Users might prioritize trust-building mechanisms as prerequisites for adoption, indicating that addressing ethical safeguards may be essential for expanding adoption beyond current users.
[…]
4. ESTIMATING FEATURE FEASIBILITY VIA PREDICTION MARKET
Table 2. Prediction market results for proposed AI features.
Feature Tool Resolved Acc. Prob. Vol. Liq.
@@ 4. ESTIMATING FEATURE FEASIBILITY VIA PREDICTION MARKET
Project-wide Context Any Tool 82% 440 290
@@
Proactive Docs Any Tool 33% 878 310
@@
Business Context Any Tool 22% 272 150
@@
Real-Time Refactor Any Tool 19% 950 370
@@
Proactive Optim. Any Tool 19% 650 310
@@
Proactive Debug Any Tool 6% 750 210
@@
Brain-computer Interface Any Tool 4% 1241 490
@@
Note: Resolved: feature launched (✓) or not (✗). Acc.: prediction correctness. Prob.: final market probability. Vol. and Liq.: participant engagement metrics.
Having identified developer needs across user groups (RQ1, RQ2), we now turn to RQ3: assessing which user needs and feature directions tool builders consider feasible and valuable and whether this reflects an accurate intuition about how the field is likely to evolve. To answer this, we conducted a prediction market “tournament” on JetBrains’ internal platform, where participants bet on whether specific AI features would be released within six months.
@@
All materials used for data gathering at this phase, as well as the full demography table and a summarized report of the prediction market, are in Supplementary materials (JetBrains Research, 2025).
4.1. METHOD
4.1.1. Materials